OpenStudy (anonymous):

how do i simplify sqrt16 + 3 sqrt 8 - 2 sqrt 2

OpenStudy (akashdeepdeb):

\[\sqrt 16 + 3 \sqrt 8 - 2 \sqrt 2 = 4 + 3 \sqrt 2*2*2 - 2 \sqrt 2 = \] \[4 + 6 \sqrt 2 - 2 \sqrt 2 = 4 + 4 \sqrt 2\] Getting me? :)

OpenStudy (anonymous):

yes thanks so much

OpenStudy (akashdeepdeb):

:)